OpenStudy (anonymous):

Integrate square root of (x^2 -4)

hartnn (hartnn):

you will get direct formula for sqrt(x^2-a^2) here http://openstudy.com/users/hartnn#/updates/50960518e4b0d0275a3ccfba or do u need to derive it ?

hartnn (hartnn):

to derive, u need to use product(uv) rule of ntegration

hartnn (hartnn):

see formula 23 in that link

hartnn (hartnn):

\(\\ \large 23. \int \sqrt{x^2-a^2}dx=\frac{x}{2}\sqrt{x^2-a^2}-\frac{a^2}{2}\ln| x+\sqrt{x^2-a^2}|+c \\\)

OpenStudy (anonymous):

So it means that what i hv 2 do is just subtitude a=2 in that rule?

hartnn (hartnn):

if you can use that standard formula, then YES. if u need to derive it, you use product rule.

OpenStudy (anonymous):

Ok,i ll try..thx ^_^

hartnn (hartnn):

if u are trying with uv rule and u don't get it, then ask. welcome :)

OpenStudy (anonymous):

Let U=√(x^2-4) dv= 1 du/dx= x/√(x^2-4) v=x dx So it become: x√(x^2-4) - ∫x^2/√(x^2-4)dx ? then do integration again for ∫x^2/√(x^2-4) dx ...stuck T.T

hartnn (hartnn):

take dv=√(x^2-4)

OpenStudy (anonymous):

Hmm...hw should i determine which 2 put u n dv?

hartnn (hartnn):

wait, which formula of uv are you using ?

hartnn (hartnn):

try this \(\\ \huge \int uv\:dx=u\int v\:dx-\int(\frac{du}{dx}\int v.dx)dx \\\) v=1 u=sqrt(....) then you will also need \(\\ \huge 19. \int \sqrt{\frac{1}{ {x^2-a^2}}}dx=\ln|x+\sqrt{x^2-a^2}|+c \\ \)

OpenStudy (anonymous):

\[x \sqrt{x ^{2}-4} - \int\limits \frac{ x ^{2} }{\sqrt{x ^{2}-4} }dx\] Then for \[\int\limits \frac{ x ^{2} }{ \sqrt{x ^{2}-4} }dx\], let u=x^2 , dv= 1/ sqrt (x^2-4) get du/dx=2x v= \[\int\limits \frac{ 1 }{ \sqrt{x ^{2}-4} }\]= \[\ln \left| x+\sqrt{x ^{2}-4} \right| +c\] then combine them by using the rule?